# Intelligence Test

### v1

@@ Abstract
How does intelligence emerge? We propose that intelligence is not a sudden gift or random occurrence, but rather a necessary trait for species to survive through Natural Selection. If a species passes the test of Natural Selection, it demonstrates the intelligence to survive in nature. Extending this perspective, we introduce Intelligence Test, a method to quantify the intelligence of any subject on any task. Like how species evolve by trial and error until they find a way to survive, Intelligence Test quantifies intelligence by the number of failed attempts before success. Fewer failures correspond to higher intelligence. When the expectation and variance of failure counts are both finite, it signals the achievement of an autonomous level of intelligence. Using Intelligence Test, we comprehensively evaluate existing artificial intelligence (AI) systems. Our results show that while AI systems achieve a level of autonomy in simple tasks, they are still far from autonomous in more complex tasks, such as vision, search, recommendation, and language. While scaling current AI technologies might help, this would come at an astronomical cost. Projections suggest that achieving autonomy for general tasks would require unimaginable $10^{26}$ parameters. To put this into perspective, loading such a massive model requires so many H100 GPUs that their total value is $4\times 10^{7}$ times that of Apple Inc.’s market value. Even with continuous hardware improvements suggested by Moore’s Law, such a parameter scale would take $70$ years to support. This staggering cost highlights the complexity of human tasks and the inadequacies of current AI technologies. To further understand this phenomenon, we conduct a theoretical analysis of Intelligence Test and its experimental results. Our findings suggest that human tasks possess a “criticality” property. As a result, autonomy requires a deep understanding of the task’s underlying mechanisms. Current AI systems, however, do not fully grasp these mechanisms and instead rely on superficial mimicry, making it difficult for them to reach an autonomous level. We believe Intelligence Test is a powerful tool. It not only guides the future development of AI but also offers profound insights into the intelligence of any species, including humans ourselves.
@@ 1 Introduction
We propose to consider the question [^1] This opening is inspired by Turing (1950). We find his framing perfectly suited to our perspective on investigating intelligence and formalize the opening in honor of his profound contributions., “What is intelligence?” This should begin with the definition of the meaning of the term “intelligence”. The definition might be framed so as to reflect so far as possible the normal use of the word, but this attitude is dangerous, If the meaning of the word “intelligence” is to be found by examining how it is commonly used, it is difficult to escape the conclusion that the answer to the question is to be sought in a statistical survey. But this is absurd. Instead of attempting such a definition, we shall replace the question with another, which is closely related to it and is expressed in relatively unambiguous words.
We turn to another question: “How does intelligence emerge?” We believe that intelligence is not an innate gift but rather a necessity shaped by Natural Selection. The diversity of life forms we see today, including humans, animals, and plants, has endured countless challenges imposed by the natural world over vast stretches of time. Take humankind as an example. We have withstood the threats of predators, survived devastating plagues, and endured two world wars. We have invented steam engines, harnessed electricity, and built the internet. None of these challenges were overcome by mere chance; none of these inventions appeared in an instant. Every step forward was the result of relentless experimentation, repeated failures, and the eventual success of those who possessed the intelligence to find a way forward. Thus, we argue that intelligence is to pass the test of Natural Selection. It drives living beings to persistently explore, seek solutions in the face of uncertainty, and ultimately prevail.
In this light, we can view Natural Selection as a test of intelligence in the domain of survival. This test challenges its subjects on their ability to endure and adapt within the natural world. The subjects must strive to solve the challenges, experimenting again and again until they succeed. If they cannot find a solution, they fail the test and thus do not survive. Those who pass this test can be considered intelligent in terms of survival. Even though the plants and animals that exist today may not have built a civilization like humans have, they have all successfully passed the test of Natural Selection. In contrast, the countless species that have gone extinct did not. In this sense, all living beings that continue to exist, including humans, possess intelligence in the task of survival.
@@
Inspired by Natural Selection, we propose Intelligence Test. Similar to how species find a way to survive through trial and error in Natural Selection, Intelligence Test quantifies intelligence by counting the number of failures a subject experiences before finding the correct solution. Fewer failures correspond to higher intelligence. When applied to practical tasks, the number of failures is a discrete random variable, and smaller expectations and variances of the failure count indicate higher intelligence. Based on the convergence of the expectations and variances of failure count, the Intelligence Test divides intelligence into three levels: Limited, Capable, and Autonomous. If both the expectation and variance diverge, the subject is at the Limited Level. At this level, the subject is comparable to blindly enumerating possible solutions. If both the expectation and variance converge, the subject reaches the Autonomous Level. At this level, the subject can stably find the correct solution with only a few trials, thereby being able to autonomously operate at an affordable cost. As we can see, the results of the Intelligence Test have clear physical meaning about the subject’s intelligence level.
@@
The Intelligence Test can be applied to any task and any species. In this paper, we are particularly interested in artificial intelligence (AI) systems. Therefore, we conduct Intelligence Test on state-of-the-art AI systems available today. The results demonstrate that a system with better modeling of the task can reach a higher level of intelligence. Current AI technologies can reach the Autonomous Level on simple tasks like handwritten digit recognition. However, they are mostly at Limited Level on more complex tasks, including vision, search, recommendation, and language. This indicates that most AI systems are at a preliminary stage: they are unable to substantially narrow down the range of possible answers and their performance is comparable to brute-force enumeration. This indicates that directly applying these AI technologies can result in very high costs and serious errors, so they cannot operate autonomously, and human supervision is essential. These findings challenge conclusions from previous studies (Biever, 2023; Aharoni et al., 2024; Mei et al., 2024), which suggest that AI has already reached a very high level of intelligence.
@@
Moreover, we empirically find that the intelligence measurement of the Intelligence Test exhibits an approximately log-linear relationship with the scale of AI systems. If we assume this relationship continues to hold, we can predict the scale required to achieve Autonomous-Level intelligence. The projection suggests that, for general language tasks, an AI system would need a parameter size of $10^{26}$ to reach the Autonomous Level. To put this scale into perspective, this is equivalent to $10^{5}$ times the total number of neurons in all of humanity’s brains combined. Loading a model of this size onto H100 GPUs would necessitate $5\times 10^{15}$ H100 cards, a cost equivalent to $4\times 10^{7}$ times the market value of Apple Inc. If hardware development continues to follow Moore’s Law, it would take 70 years of progress to support the development of such a large model. These results suggest that attempting to solve human tasks with current AI technology is extremely difficult, if not impossible.
@@
Why is the Autonomous Level so difficult to achieve for current AI systems? We conduct a theoretical analysis and demonstrate that the root cause lies in the complexity of human tasks and the inadequacies of current AI technologies. Specifically, we leverage self-organized criticality (SOC) (Bak, 2013) to analyze Intelligence Test. Results suggest that many human tasks exhibit a criticality property: even slight changes in the environment require entirely different responses. To successfully operate these tasks, it is important to fully understand their mechanisms. However, current AI systems do not fully grasp this complex mechanism and instead leverage superficial imitation: They memorize answers to some questions and attempt to solve new questions through exploration. Although scaling AI systems can make the exploration more effective and improve imitation performance, a lack of a full understanding of the underlying mechanism results in unimaginable costs.
@@
The structure of this paper is as follows. In Section 2, we review related works on intelligence evaluation to provide a broad context for our method. In Section 3, we present Intelligence Test and show how it measures intelligence and categorizes it into three levels. In Section 4, we extensively evaluate existing AI systems using Intelligence Test. Section 5 empirically explores how scaling improves Intelligence. In Section 6, we provide a theoretical analysis of Intelligence Test to gain a deep understanding of the nature of human tasks and current AI. Finally, in Section 7, we conclude the paper and outline potential directions for future research.
@@ 2 Related Work
We observe that almost all previous works attempted to define intelligence by determining which human-like tasks a machine must accomplish to be considered intelligent. However, there is significant variation in the choice of the tasks, as different researchers hold different perspectives on what constitutes intelligence. We can see that these approach approaches are inherently subjective, which manifests in three key ways:
  • Subjective (Human-Centric) View of Intelligence: Many of these tests utilize human intelligence as an upper bound for AI and evaluate whether AI can approach this bound. For example, Imitation Game (Turing, 1950) evaluates a machine’s ability to replicate human behavior; OpenAI (2018) defines intelligence as outperforming humans at economically valuable work; Morris et al. (2024) defines intelligence level by how many humans AI can outperform. Nevertheless, humans are not always the most intelligent in every task and should not serve as an upper bound. Assuming an upper bound makes the methods no longer applicable once AI surpasses this bound.
@@
In contrast, our proposed Intelligence Test has a clear physical meaning and a well-defined statistical basis. It is inherently an objective way to evaluate intelligence. Before we further elaborate on the differences between Intelligence Test and related studies, we will first introduce Intelligence Test in Section 3 and then continue the comparison in Section 3.4.
@@ 3 Methodology
The origin of species, including humans, is Natural Selection. Nature has relentlessly filtered out the unintelligent, allowing only those with intelligence to survive. Inspired by this, we propose Intelligence Test, a universal framework to quantify the intelligence of any subject in any task. It builds upon the core concept of Natural Selection to test how well a subject can autonomously explore and find solutions.
In the following subsections, we first revisit Natural Selection and formalize it as a Survival Game. Then, we extend Survival Game as Intelligence Test to quantify intelligence at any task. Next, we interpret the results of Intelligence Test into three levels of intelligence. Finally, we discuss how Intelligence Test differs from previous studies.
3.1 Survival Game
The process of Natural Selection is extraordinarily complex, involving competition between species, genetic mutations, etc. Rather than delving into these intricate details, we simplify Natural Selection into a conceptual framework which we call the Survival Game.
@@ 3.1 Survival Game
Despite the simplification, Survival Game captures the essence of Natural Selection. Throughout history, nature has posed countless challenges to humankind. When asked how to survive predators, the intelligent among us answered fire and tools. When faced with the threat of starvation, the intelligent among us developed agriculture. When confronted with disease, the intelligent among us advanced medicine. Civilization itself has been forged through these relentless Survival Games. If it were not always for someone to step up and pass the Survival Game, modern society as we know it would not exist.
Based on the description of the Survival Game above, we will now translate it into mathematical terms to make it clearer.
Definition 3.2 (Survival Game in Math Language)
We can see that the number of failures, $X$, is a direct measure of a species’ survival intelligence. The smaller the value of $X$, the less effort the species needs to solve problems. Inspired by this, our proposed Intelligence Test will similarly measure intelligence.
3.2 Measuring Intelligence on Any Task
Survival Game directly measures the intelligence of species on the survival task. Building on this, we introduce Intelligence Test, which extends the Survival Game to any task and measures any subject’s intelligence. To achieve this, Intelligence Test models failure counts as a discrete random variable and uses statistical metrics for evaluation. The modifications are two-fold:
  • Modeling Failure Count as a Discrete Random Variable: One task can involve numerous variations, and the conclusions of the Survival Game may be very different across these variations. For instance, consider testing a subject’s ability to solve mathematical problems. A small change in the numbers or the context of the problem could lead to a significant shift in the subject’s failure counts. Similarly, when a task is classifying images, different pictures can result in substantial fluctuations in performance. Therefore, the variability within the task can cause the conclusions to be unstable. To account for this variability, Intelligence Test models failure count as a discrete random variable, which allows us to handle the variations across task variants effectively.
  • Statistical Criteria for Evaluation: Population size $N$ serves as a threshold value in the Survival Game. It directly affects the conclusion. The larger the value of $N$, the more attempts are available to the subject, and consequently, the higher the likelihood of success. Yet, for tasks other than survival, the notion of what constitutes an “appropriate” $N$ can vary from one researcher to another. This variability in determining an appropriate N leads to inconsistencies in the conclusions. Therefore, Intelligence Test does not use a pre-defined threshold for measurement. It quantifies intelligence as the distribution of failure count. A lower probability of a large failure count suggests higher intelligence.
With these statistical improvements, we formally define Intelligence Test as follows:
Definition 3.3 (Intelligence Test)
@@ 3.2 Measuring Intelligence on Any Task
Let $X$ be a discrete random variable representing the number of a subject’s failure attempts before providing a correct solution for a task. Then, $X$ serves as the measure of this subject’s intelligence on the task. Smaller expectations and variances of $X$ correspond to higher intelligence.
@@
Smaller expectations and variances indicate that the subject can achieve success with fewer failures and thus is more intelligent. This definition allows us to assess intelligence in any given task and for any species. We can choose to evaluate intelligence in narrow tasks such as answering domain-specific questions, or we can test a subject across diverse and complex tasks to determine whether it exhibits general intelligence, such as memorizing every information on the Internet. Besides, this framework is not restricted to any particular species. It can be applied to humans, other biological species, or artificial intelligence systems. The measurement of Intelligence Test has a clear physical meaning: it signifies how well a subject can reliably find solutions for a given task on its own.
Note that Intelligence Test requires to determine whether each attempt made by a subject is correct. Yet verifying correctness for every attempt can be expensive in some tasks. Consider a task where the test subject is to prove a mathematical theorem. The subject provides proof with each attempt. However, for complex mathematical theorems, the proofs can be very long and intricate, and the cost of verifying the correctness of each proof is extremely high. In such cases, directly applying Intelligence Test may make the process prohibitively expensive.
To practically apply Intelligence Test when verifying every attempt is costly, we propose a variation named Intelligence Test with References. It avoids the need for direct correctness verification while keeping the core of Intelligence Test. In those tasks where the cost of correctness verification is high, it uses a reference answer and measures the number of failed attempts before producing the reference answer. The validity of this method is supported by the following theorem:
Theorem 3.4 (Intelligence Test with References)
Let $X^{*}$ be a discrete random variable representing the number of unsuccessful attempts before discovering a predefined reference answer in a task. $X^{*}$ is an upper bound estimate of $X$ in Intelligence Test.
Proof: Note that the reference answer is also a valid solution. The failure counts to reach one specific solution must be greater than or equal to the failure counts for any solution.
This approach eliminates the need for verifying every attempt and instead examines failure counts until reaching a known reference answer. For example, when testing whether a subject can prove mathematical theorems, this test checks whether it can produce a known proof, rather than evaluating whether each of its outputs constitutes a valid new proof. It is a low-cost realization of Intelligence Test. In practice, whether to use this variation depends on the cost of verification.

### v2

# Evaluating Intelligence via Trial and Error

@@ Abstract
How does intelligence emerge? We propose that intelligence is not a sudden gift or random occurrence, but rather a crucial trait for species to survive through Natural Selection. Natural Selection requires a species to find solutions within a limited number of trial-and-error attempts. Building on this idea, we introduce Survival Game as a framework to evaluate intelligence based on the number of failed attempts in a trial-and-error process. Fewer failures indicate higher intelligence. When the expectation and variance of failure counts are both finite, it signals the ability to consistently find solutions to new challenges, which we define as the Autonomous Level of intelligence. Using Survival Game, we comprehensively evaluate existing artificial intelligence (AI) systems. Our results show that while AI systems achieve the Autonomous Level in simple tasks, they are still far from it in more complex tasks, such as vision, search, recommendation, and language. While scaling current AI technologies might help, this would come at an astronomical cost. Projections suggest that achieving the Autonomous Level for general tasks would require $10^{26}$ parameters. To put this into perspective, loading such a massive model requires so many H100 GPUs that their total value is $4\times 10^{7}$ times that of Apple Inc.’s market value. Even with Moore’s Law, supporting such a parameter scale would take $70$ years. This staggering cost highlights the complexity of human tasks and the inadequacies of current AI technologies. To further investigate this phenomenon, we conduct a theoretical analysis of Survival Game and its experimental results. Our findings suggest that human tasks possess a “criticality” property. As a result, Autonomous Level requires a deep understanding of the task’s underlying mechanisms. Current AI systems, however, do not fully grasp these mechanisms and instead rely on superficial mimicry, making it difficult for them to reach an autonomous level. We believe Survival Game can not only guide the future development of AI but also offer profound insights into human intelligence.
@@ 1 Introduction
How does intelligence emerge? We believe that intelligence is not an innate gift but rather a necessity shaped by Natural Selection. The diversity of life forms we see today, including humans, animals, and plants, has addressed countless challenges imposed by the natural world. Natural Selection can be viewed as a trial-and-error process, where subjects shall persistently explore, seek solutions in the face of uncertainty, and ultimately prevail. The subjects must strive to solve the challenges, experimenting again and again until they succeed. If they cannot find a solution, they fail the test and thus do not survive.
@@
Inspired by Natural Selection, we propose Survival Game to evaluate intelligence. Similar to how species find a way to survive through trial and error in Natural Selection, Survival Game evaluates intelligence by counting the number of failures before finding correct solutions in a trial-and-error process. Fewer failures correspond to higher intelligence. The number of failures is a discrete random variable, and smaller expectations and variances of the failure count indicate higher intelligence. If expectations and variances are infinite, the subjects can never find the correct solutions and thus do not survive in the game. Based on the convergence of the expectations and variances, Survival Game divides intelligence into three levels: Limited, Capable, and Autonomous. If both the expectation and variance diverge, the subject is at the Limited Level. At this level, the subject is comparable to blindly enumerating possible solutions. If both the expectation and variance converge, the subject reaches the Autonomous Level. At this level, the subject can stably find the correct solution with only a few trials, thereby being able to autonomously operate at an affordable cost. As we can see, the results of the Survival Game have clear physical meaning about the subject’s intelligence level.
@@
The Survival Game can be applied to any task and any species. In this paper, we are particularly interested in artificial intelligence (AI) systems. Therefore, we conduct Survival Game on state-of-the-art AI systems available today. The results demonstrate that a system with better modeling of the task can reach a higher level of intelligence. Current AI technologies can reach the Autonomous Level on simple tasks like handwritten digit recognition. However, they are mostly at Limited Level on more complex tasks, including vision, search, recommendation, and language. This indicates that most AI systems are at a preliminary stage: they are unable to substantially narrow down the range of possible answers and their performance is comparable to brute-force enumeration. This indicates that directly applying these AI technologies can result in very high costs and serious errors, so they cannot operate autonomously, and human supervision is essential. These findings challenge conclusions from previous studies (Biever, 2023; Aharoni et al., 2024; Mei et al., 2024), which suggest that AI has already reached a very high level of intelligence.
Figure 1: Experimental Results of Survival Game in General Domain. Left: Results suggest that larger models achieve better performance. Right: Results suggest that achieving Autonomous-Level Intelligence requires an unimaginable parameter scale.
@@
In Survival Game, the intelligence score exhibits a log-linear relationship with the scale of AI systems. If we assume this relationship continues to hold, we can predict the scale required to achieve Autonomous-Level intelligence, as shown in Figure 1. The projection suggests that, for general language tasks, an AI system would need a parameter size of $10^{26}$ to reach the Autonomous Level. To put this scale into perspective, this is equivalent to $10^{5}$ times the total number of neurons in all of humanity’s brains combined. Loading a model of this size onto H100 GPUs would necessitate $5\times 10^{15}$ H100 cards, a cost equivalent to $4\times 10^{7}$ times the market value of Apple Inc. If hardware development continues to follow Moore’s Law, it would take 70 years of progress to support the development of such a large model. These results suggest that attempting to solve human tasks with current AI technology is extremely difficult, if not impossible.
@@
Why is the Autonomous Level so difficult to achieve for current AI systems? We conduct a theoretical analysis and demonstrate that the root cause lies in the complexity of human tasks and the inadequacies of current AI technologies. Specifically, we leverage self-organized criticality (SOC) (Bak, 2013) to analyze Survival Game. Results suggest that many human tasks exhibit a criticality property: even slight changes in the environment require entirely different responses. To successfully operate these tasks, it is important to fully understand their mechanisms. However, current AI systems do not fully grasp this complex mechanism and instead leverage superficial imitation: They memorize answers to some questions and attempt to solve new questions through exploration. Although scaling AI systems can make the exploration more effective and improve imitation performance, a lack of a full understanding of the underlying mechanism results in unimaginable costs.
@@
The structure of this paper is as follows. In Section 2, we review related works on intelligence evaluation to provide a broad context for our method. In Section 3, we present Survival Game and show how it measures intelligence and categorizes it into three levels. In Section 4, we extensively evaluate existing AI systems using Survival Game. Section 5 empirically explores how scaling improves Intelligence. In Section 6, we provide a theoretical analysis of Survival Game to gain a deep understanding of the nature of human tasks and current AI. Finally, in Section 7, we conclude the paper and outline potential directions for future research.
@@ 2 Related Work
We observe that almost all previous works attempted to define intelligence by determining which human-like tasks that a machine must accomplish in order to be considered intelligent. However, there is significant variation in the choice of the tasks, as different researchers hold different perspectives on what constitutes intelligence. We can see that these approach approaches are inherently subjective, which manifests in three key ways:
  • Subjective (Human-Centric) View of Intelligence: Many of these tests utilize human intelligence as an upper bound for AI and evaluate whether AI can approach this bound. For example, Imitation Game (Turing, 1950) evaluates a machine’s ability to replicate human behavior; OpenAI (2018) defines intelligence as outperforming humans at economically valuable work; Morris et al. (2024) defines intelligence level by how many humans AI can outperform. Nevertheless, if AI surpasses humans in certain tasks, these evaluation methods are no longer applicable.
@@
In contrast, our proposed Survival Game has a clear physical meaning and a well-defined statistical basis. It is inherently an objective way to evaluate intelligence. Before we further elaborate on the differences between Survival Game and related studies, we will first introduce Survival Game in Section 3 and then continue the comparison in Section 3.5.
@@ 3 Methodology
We propose Survival Game, a framework to evaluate intelligence via a trial-and-error process. Its core concept is to test how well a subject can autonomously explore and find solutions. In the following subsections, we first revisit Natural Selection and formalize it as a Survival Game. Then, we extend Survival Game as Survival Game to quantify intelligence at any task. Next, we interpret the results of Survival Game into three levels of intelligence. Furthermore, we propose an approximation method to apply Survival Game in costly tasks. Finally, we discuss how Survival Game differs from previous studies.
3.1 Natural Selection as a Trial-and-Error Test
Natural Selection is an intuitive way to test intelligence. If a subject passes Natural Selection, it signifies that this subject possesses the intelligence to operate autonomously and can sustain itself without external guidance. The process of Natural Selection is extraordinarily complex, involving competition between species, genetic mutations, etc. Rather than delving into these intricate details, we simplify Natural Selection into a trial-and-error test as follows:
Definition 3.1
Despite the simplification, this trial-and-error test captures the essence of Natural Selection. Throughout history, nature has posed countless challenges to humankind. When asked how to survive predators, the intelligent among us answered fire and tools. When faced with the threat of starvation, the intelligent among us developed agriculture. When confronted with disease, the intelligent among us advanced medicine. Civilization itself has been forged through these relentless trials and errors.
Based on the description of the above trial-and-error process, we can translate it into mathematical terms to make it clearer.
Definition 3.2
We can see that the number of failures, $X$, is a direct measure of a species’ survival intelligence. The smaller the value of $X$, the less effort the species needs to solve problems. Inspired by this, our proposed Survival Game will similarly measure intelligence.
3.2 Measuring Intelligence with Survival Game
Based on the trial-and-error process in Natural Selection, we introduce Survival Game, which evaluates intelligence by the number of failure attempts in this process. To ensure a robust evaluation result, Survival Game models failure counts as a discrete random variable and uses statistical metrics for evaluation. The modifications are two-fold:
  • Modeling Failure Count as a Discrete Random Variable: One task can involve numerous variations, and failure counts may be very different across these variations. For instance, consider testing a subject’s ability to solve mathematical problems. A small change in the numbers or the context of the problem could lead to a significant shift in the subject’s failure counts. Similarly, when a task is classifying images, different pictures can result in substantial fluctuations in performance. Therefore, the variability within the task can cause the results to be unstable. To account for this variability, Survival Game models failure count as a discrete random variable, which allows us to handle the variations across task variants effectively.
  • Statistical Criteria for Evaluation: Population size $N$ serves as a threshold value in the trial-and-error test for Natural Selection. It directly affects the conclusion. The larger the value of $N$, the more attempts are available to the subject, and consequently, the higher the likelihood of success. Yet, for tasks other than survival, the notion of what constitutes an “appropriate” $N$ can vary from one researcher to another. This variability in determining an appropriate N leads to inconsistencies in the conclusions. Therefore, Survival Game does not use a pre-defined threshold for measurement. It quantifies intelligence as the distribution of failure count. A lower probability of a large failure count suggests higher intelligence.
With these statistical improvements, we formally define Survival Game as follows:
@@ 3.2 Measuring Intelligence with Survival Game
Let a subject perform a certain task through continuous trial and error until finding the correct solution. $X$ is a random variable representing the number of failures before the subject finds the correct solution. Then, $X$ serves as the measure of this subject’s intelligence on the task. Smaller expectations and variances of $X$ correspond to higher intelligence.
@@
Smaller expectations and variances indicate that the subject can achieve success with fewer failures and thus is more intelligent. This definition allows us to assess intelligence in any given task. We can choose to evaluate intelligence in narrow tasks such as answering domain-specific questions, or we can test a subject across diverse and complex tasks to determine whether it exhibits general intelligence, such as memorizing every information on the Internet. The measurement of Survival Game has a clear physical meaning: it signifies how well a subject can reliably find solutions for a given task on its own.
It is worth noting that Survival Game assumes that subjects must keep trying until they succeed even if the cases are very difficult. For easy cases where subjects can answer correctly without trial and error, the contribution to the failure count is zero. We can see that Survival Game essentially ignores easy cases where subjects can answer correctly right away and instead focuses on difficult cases that require repeated trials and errors. For example, in image classification, Survival Game focuses on images that the AI model initially misclassifies. It examines how many trial-and-error attempts are needed before achieving the correct classification. This emphasis on trial and error differentiates the Survival Game from existing evaluation methods based on accuracy. In real-world applications, if a task is highly sensitive to errors, such as high-risk decision-making scenarios like autonomous driving, Survival Game provides a better reflection of whether an AI model can be trusted. Additionally, in highly intellectual tasks that require AI to go through trial and error to find a solution, such as proving mathematical theorems or optimizing agent workflows, Survival Game metric directly corresponds to computational cost and shall better reflect AI’s applicability.